A stone is thrown vertically upwards with a speed of 20.0 m / s. (a) What is its speed when it reaches 18 meters? (b) What maximum height does it reach?
Answer #1

3 Udd > SIUNIWUJO a) v at 1 1 1 - h=18m1 dg M = +20m/s (initial velocity) pu=20ms a=-9.8 m/s2=g ch=displacement = +18m V = sp